OpenStudy (anonymous):

What is the equation of the line that passes through the point (2, 5) and has a slope of 2? y = 2x - 3 y = 2x - 8 y = 2x + 1 y = 2x + 3

OpenStudy (jdoe0001):

http://www.teachhandson.com/calculationb.gif

OpenStudy (ivettef365):

the slope is 2, use formula y=mx+b to find b 5 = 2(2) + b 5-4 = b then the equation = y = 2x + 1

OpenStudy (ivettef365):

that is the correct answer